if tanA=1/2, find the value of cosA/sinA+sinA/1+cosA

Grade:10

1 Answers

SAI SARDAR
8 years ago
(5+2√5)÷(2+√5).if tanA=1/2, find the value of cosA/sinA+sinA/1+cosA.then answer is
because tanA=1/2,so sinA=1/root 5 and cos A =2/root 5.